Solve for v, when E=1/2mv to the 2nd power.
Answer by stanbon(75887)

Solve for v, when E=1/2mv to the 2nd power.
-------------
E = (1/2)mv^2
-----
2E = mv^2
v^2 = [2E/m]
v = sqrt[2E/m]
------------------
Cheers,
Stan H.
----------------